OpenStudy (anonymous):

If f(x)=3*square root of (2x-1) - 3*square root of (x-4), find x so that f(x) =6

OpenStudy (raden):

haiii.. are u still there

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

can you solve this problem for me

OpenStudy (raden):

sure, i will explain it now... listen to me :)

OpenStudy (raden):

f(x) = 6 3 sqrt(2x-1) - 3 sqrt(x-4) = 6 simplify, first. divided by 3 to both sides, giving us : sqrt(2x-1) - sqrt(x-4) = 2

OpenStudy (raden):

move side - sqrt(x-4) to right sides, it can be : sqrt(2x-1) = sqrt(x-4) + 2

OpenStudy (anonymous):

what is the SUM of these two x values

OpenStudy (anonymous):

There are two such x values. What is the SUM of these two x values?

OpenStudy (raden):

now, squaring to both sides :